# Formatting Instructions for ICLR 2025 Conference Submissions Residual Connections Harm Generative Representation Learning

### v3

@@ 4.3 Ablation Studies on ImageNet-100
| 0.5 | 0.6 | 0.7 | 0.8 | 0.9 | 1.0 | |
| ViT-B/16 | 82.3 | 83.6 | 81.8 | 79.8 | 79.2 | 76.5 |
| ViT-S/16 | 78.6 | 78.5 | 78.1 | 75.2 | 73.5 | 69.2 |
Decay rate . The only parameter of our framework is , the minimum scaling factor applied to the identity shortcut at the final layer. In Table 2,we show linear probing scores for varying values of . We observe that must be sufficiently small to regularize the flow of information through the residual connection effectively. A reasonably small prevents the deeper layers of the encoder from relying heavily on the residual connections, allowing for more abstract representations in the bottleneck. This yields up to a 7.1% improvement over the ViT-B/16 baseline () in linear probing accuracy on ImageNet-100. On the other hand, if is too small, for example, for ViT-B/16, we observe that the training becomes unstable.
Architecture size. In Table 2, we also train both the ViT-B/16 and the smaller ViT-S/16 backbone using varying . Our framework is especially effective on the smaller architecture, increasing linear probing performance by 9.4%, compared to the baseline setting (standard Residual Connection at ). This is consistent with the observation that larger models generalize better (Huh et al., 2021) than smaller models. Our method is able to significantly improve the smaller ViT-S/16 and slightly close the gap between the differently-sized architectures.
@@
Skip connections. Another critical design choice in our network is to include skip connections that are not in the original MAE. As discussed in Section 3.2, if the MAE does not use skip connections, the bottleneck layer must preserve all information to reconstruct the input image accurately. This is opposed to learn abstract representations at bottleneck. These contrary effects significantly degrade the representation learned by the model, leading to a 22.1% drop in the linear probing score, as we report in Table LABEL:tab:skip_connections.
@@
Different decay schema. We also explore decay schema, with results summarized in Table LABEL:tab:decay_schema: (1) Scaling both branches of the residual blocks simultaneously by applying a constant factor, , to both and . (2) Scaling only using the same constant factor, . (3) Applying our proposed schema exclusively to either the attention or MLP branch.
@@
Among these, (2) shows no significant improvement over the baseline, while (1) yields some improvement but still underperforms compared to our approach. By analyzing (1) and (2), we demonstrate that the representation gains are due to down-weighting the skip connection branch. Notably, recent diffusion models (Karras et al., 2018; Song et al., 2021b; Karras et al., 2020; 2022) have employed (1) in their designs. However, applying decay only to the MLP or attention branch reduces the overall decaying effect across the network, resulting in lower performance compared to our schema, which achieves the best performance among the tested designs.
Configurations UNet Accuracy No 61.5 Yes 83.6*
Configurations Decay Block Accuracy - 76.5 MLP & Atten. 76.9 MLP & Atten. 82.6 Atten. 0.6 79.3 MLP 0.6 80.6 MLP & Atten. 0.6 83.6*
@@ 6 Discussion on feature rank
where denotes the normalized singular value. Intuitively, this measure is small when a few singular values dominate and large when singular values are evenly spread, hence giving a good continuous approximation for matrix rank. In the following subsections, to compute the effective rank, we use the singular values from the covariance matrix of the last-layer features, where denotes the covariance of the learned class tokens for the and samples.
T
[…]
| coefficient | |
| 200 | -0.51 |
| 300 | -0.56 |
| 400 | -0.57 |
| 500 | -0.58 |
| 600 | -0.52 |
| 700 | -0.44 |
| 800 | -0.42 |
Inspired by Huh et al. (2021), we conjecture that the improvement to feature learning capability of our method can mainly be attributed to the decayed identity shortcuts promoting low-rank features at the bottleneck. We measure the training dynamics of the models presented in Table 2 (MAEs trained on ImageNet-100) in terms of accuracy and the effective rank. In Figure 6(c), we quantify the correlation between effective feature rank and probing accuracy using Pearson correlation coefficient to validate the hypothesis.
In Figures 6(a) and 6(b), we present the training dynamics of our model, highlighting effective rank and linear probing accuracy for different values of . During the early epochs, models with lower tend to exhibit both lower effective rank and higher probing accuracy, supporting our hypothesis. As training progresses, the correlations between and effective rank become less precise. We suspect this is due to the training dynamics, where models with lower experience faster growth in feature rank, reflecting the model’s effort to compensate for the decay factors. Despite this, we can still conclude that lower results in lower feature rank and better probing accuracy compared to higher .
@@ Appendix A Training and Evaluation Details
Our training configurations primarily followed the guidelines established by He et al. (2022). In the ImageNet-1K experiment, our model was trained for 800 epochs, utilizing the AdamW Loshchilov & Hutter (2019) optimizer with a constant weight decay of 5e-2 for a batch size of 1024. We set the maximum learning rate to 6e-4. Initially, the learning rate started at 0 and linearly increased to its maximum over the first 40 epochs, after which it followed a cosine schedule to gradually decrease to zero by the end of the training period. It is worth noting that the learning rate per sample, or effective learning rate, in our setup matched that of He et al. (2022), although our maximum learning rate was set lower due to our batch size being a quarter of theirs. We applied random resizing, cropping, and horizontal flipping during training as part of our augmentation scheme. To enhance the quality of the learned representations in most experiments, we employed the normalized pixel loss, as proposed by He et al. (2022). In the ImageNet-100 experiment, we employed the identical training configuration used in the ImageNet-1K experiments. We train our model with 4 NVIDIA A40 GPUs and a completed trianing usually takes 20 hours on ImageNet-100 and 200 hours on ImageNet-1k.
@@ A.3 Modified Architecture
We present a visualization of our UNet transformer design, as outlined in Section 3.2, in Fig. 7. It’s important to note that decayed identity shortcuts are exclusively implemented within the encoder block. Additionally, we establish skip connections from alternating blocks in the encoder to the decoder, following the UNet Ronneberger et al. (2015) architecture’s design principles.
@@ Appendix B Further experiments
B.1 Further Ablation of Maximum Decay Rate on ImageNet-1K
| Dataset | 0.5 | 0.6 | 0.7 | 0.8 | 0.9 | 1.0 | |
| ImageNet-1K | ViT-B/16 | 69.8 | 72.7 | 68.9 | - | - | - |
| ImageNet-100 | ViT-B/16 | 82.3 | 83.6 | 81.8 | 79.8 | 79.2 | 76.5 |
| ImageNet-100 | ViT-S/16 | 78.6 | 78.5 | 78.1 | 75.2 | 73.5 | 69.2 |
We present the results of ablating the choices of on ImageNet-1K dataset in Table.5. From the table, we show that the optimal for ViT-B/16 on ImageNet-100 matches the optimal one for ImageNet-1K, while a lower is preferred for a smaller architecture ViT-S/16.
B.2 Reconstruction quality.
@@ B.1 Further Ablation of Maximum Decay Rate on ImageNet-1K
B.3 Abstraction and Low-rank in the Supervised Setting

### v4

@@ 4.2 Ablation Studies on ImageNet-100
Decay rate . The only parameter of our framework is , the minimum scaling factor applied to the final layer. In Table 2,we show linear probing scores for varying values of . We observe that works well for most cases. If is too small, for example, , we observe that the training becomes unstable.
Architecture size. In Table 2, we also run experiments over multiple architecture choices. We notice encoder depth rather than feature dimension that influences the optimal choices of and deeper models requires larger . We attribute this phenomenon to the cumulative decaying effect of the final layer, quantified as . Heavy decaying would harm the optimization and selecting such that [1e-3, 1e-2) yields the best performance.
@@
Skip connections. Another critical design choice in our network is to include skip connections that are not in the original MAE. As discussed in Section 3.2, if the MAE does not use skip connections, the bottleneck layer must preserve all information to reconstruct the input image accurately. This is opposed to learn abstract representations at bottleneck. These contrary effects significantly degrade the representation learned by the model, leading to a 22.1% drop in the linear probing score, as we report in Table 3(a).
Scheduler. We use linear scheduler as a default choice. In table 3(a), we also experiment with a cosine scheduler but find it leads to worse performance for . Besides using a prefixed scheduler, we also experiment with a learnable , which resembles the setup of highway network (Srivastava et al., 2015a). We show the learnable at each layer in Table5, appendix. From the table, we don’t find consistent patterns over network depth and the performance is worse than our predefined scheduler.
@@
Different decay schema. We also explore decay schema, with results summarized in Table 3(b): (1) Scaling both branches of the residual blocks simultaneously by applying a constant factor, , to both and . (2) Scaling only using the same constant factor, . (3) Applying our proposed schema exclusively to either the attention or MLP branch.
@@
Among these, (2) shows no significant improvement over the baseline, while (1) yields some improvement but still underperforms compared to our approach. By analyzing (1) and (2), we demonstrate that the representation gains are due to down-weighting the skip connection branch. Notably, recent diffusion models (Karras et al., 2018; Song et al., 2021b) have employed (1) in their smaller convolutional neural network but don’t provide systemic analysis. However, applying decay only to the MLP or attention branch reduces the overall decaying effect across the network, resulting in lower performance compared to our schema, which achieves the best performance among the tested designs.
4.3 Embdding analysis
We qualitatively evaluate the feature learning in Figure 3 and we adopt the pixel-wise embedding approaches proposed by Zhang et al. (2024) to group the representations from the last layer of the encoder into a lower dimensional space. We use their default hyperparameters to cluster representations across COCO validation set and render the top 3 eigenvectors as RGB channels of images. From the visualization, ours learns abstract representation and the object from the same categories have similar color, indicating a global consistent semantic grouping. The baseline MAE, on the other hand, doesn’t show clearly global semantic patterns.
We also benchmark the clustering quantitatively, following the postprocessing protocol (Zhang et al., 2024) to produced unsupervised semantic segmentation and report the results as the mean intersection of union (mIoU). Ours (10.41 mIoU) achieve 6.31 mIoU improvement over baseline (4.10 mIoU), which support the visual comparison.
@@ 6 Discussion on feature rank
where denotes the normalized singular value. Intuitively, is small when a few singular values dominate and large when singular values are evenly spread, hence giving a good continuous approximation for matrix rank. In the following subsections, we use the singular values from the covariance matrix of the last-layer features to compute , where denotes the covariance of the learned class tokens for the and samples.
Inspired by Huh et al. (2021), we conjecture that the improvement to feature learning capability of our method can mainly be attributed to the decayed identity shortcuts promoting low-rank features at the bottleneck. In Figures 4(a) and 4(b), we measure the training dynamics of the models (feature dimension 768 and encoder depth 12) in terms of accuracy and the effective rank, for different values of
During the early epochs, models with lower tend to exhibit both lower effective rank and higher probing accuracy, supporting our hypothesis. As training progresses, the correlations between and effective rank become less precise. We suspect this is due model’s effort to compensate for the decay factors with large decaying factor. Despite this, we can still conclude that lower [0.5-0.6] results in lower feature rank and better probing accuracy compared to higher = [0.7-1.0].
Compatibility with contrastive learning frameworks. Despite substantial improvement of applying our decayed identity shortcuts in generative model, we note that our approach does not easily extend to contrastive learning frameworks, where the low rank inductive bias conflicts with the training objectives, e.g., MoCov3 (Chen et al., 2021b) include an universal repulsive term in the denominator to increase the feature rank. Rankme (Garrido et al., 2023) confirms this by showing conservative learning models with higher feature ranks yield better results.
@@ 7 Conclusion
8 Impact Statements
This paper presents work whose goal is to advance the field of generative representation learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Training and Evaluation Details
Our training configurations primarily followed the guidelines established by He et al. (2022). In the ImageNet-1K experiment, our model was trained for 800 epochs, utilizing the AdamW (Loshchilov & Hutter, 2019) optimizer with a constant weight decay of 5e-2 for a batch size of 1024. We set the maximum learning rate to 6e-4. Initially, the learning rate started at 0 and linearly increased to its maximum over the first 40 epochs, after which it followed a cosine schedule to gradually decrease to zero by the end of the training period. It is worth noting that the learning rate per sample, or effective learning rate, in our setup matched that of He et al. (2022), although our maximum learning rate was set lower due to our batch size being a quarter of theirs. We applied random resizing, cropping, and horizontal flipping during training as part of our augmentation scheme. To enhance the quality of the learned representations in most experiments, we employed the normalized pixel loss, as proposed by (He et al., 2022). In the ImageNet-100 experiment, we employed the identical training configuration used in the ImageNet-1K experiments. We train our model with 4 NVIDIA A40 GPUs and a completed trianing usually takes 20 hours on ImageNet-100 and 200 hours on ImageNet-1k.
@@ A.3 Modified Architecture
We present a visualization of our UNet transformer design, as outlined in Section 3.2, in Fig. 5. It’s important to note that decayed identity shortcuts are exclusively implemented within the encoder block. Additionally, we establish skip connections from alternating blocks in the encoder to the decoder, following the UNet (Ronneberger et al., 2015) architecture’s design principles.
A.4 Learnable over network layers
| Layer Index | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 11 | 12 |
| Attention | 0.993 | 0.947 | 0.982 | 0.766 | 0.992 | 0.795 | 0.988 | 0.849 | 0.998 | 0.723 | 0.811 | 0.488 |
| FFN | 0.989 | 0.926 | 0.961 | 0.620 | 0.961 | 0.442 | 0.711 | 0.322 | 0.810 | 0.475 | 0.637 | 0.353 |
In the ablation study of learnable , we apply no additional regularization beyond standard weight decay to the model parameters. Table 5 presents the values for each layer. The results do not reveal any meaningful pattern across network depth.
@@ Appendix B Further experiments
B.1 Reconstruction quality.
@@
B.2 Abstraction and Low-rank in the Supervised Setting